OpenStudy (anonymous):

I have 4 questions that have to do with logs. If someone could help, that'd be great! 1. log6 (10)= x Log6 (7)= y Log6 (4) = z Find log6 (1/9) 2. Log7 (3) = x Log7 (8) = y Log7 (10) = z Find log7 (15/32) 3. Log8 (10)= p Log8 (12) = q Log8 (11) = r Find log8 (5/256) 4. Log8 (12) = p Log8 (5) = q Log8 (9) = r Find log8 (32/27)

OpenStudy (anonymous):

https://www.desmos.com/calculator This should help! :)

OpenStudy (anonymous):

Thank you...Only thing is that I can't really use set values for these logs, so...

OpenStudy (mertsj):

Could you please post the directions for these problems?

OpenStudy (anonymous):

Here are the directions: use the properties of logarithms and the values below to find the logarithm indicated, do not use a calculator to evaluate the logs.

OpenStudy (anonymous):

wow that is a tricky one

OpenStudy (anonymous):

the idea is to figure out how to write \(\frac{1}{9}\) as a combination of \(4,7,10\) and \(6\)

OpenStudy (anonymous):

only operations allowed are multiply, divide, raise to powers, and take roots to get \(\frac{1}{9}\) you can use \[\frac{1}{9}=\left(\frac{1}{3}\right)^2=\left(\frac{\sqrt{4}}{6}\right)^2\]

OpenStudy (anonymous):

so \[\log(\frac{1}{9})=\log\left(\frac{\sqrt{4}}{6}\right)^2\]

OpenStudy (anonymous):

guess i lost her

OpenStudy (anonymous):

actually \[\log(\frac{4}{6})^2\] would work better

OpenStudy (anonymous):

Yeah ; w ; I'm still pretty stumped..

OpenStudy (anonymous):

Huhu, thanks :)